OpenStudy (pphalke):

A lab technician mixed a 530 ml solution of water and alcohol. If 3% of the solution is alcohol, how many milliliters of alcohol were used and how many milliliters of water were used?

OpenStudy (abdullahm):

3% = 3/100 3/100 * 530 = milliliters of alcohol 100% - 3% = 97% = 97/100 97/100 * 530 = milliliters of water

OpenStudy (pphalke):

woah! wait those both answers are the same but that doesnt seem right to me @AbdullahM

OpenStudy (abdullahm):

Both answers are not the same.

OpenStudy (benthicflare):

@pphlake I thought you posted this before?

OpenStudy (pphalke):

no i just posted it like 9 min ago where did u see this? maybe it was a while back

OpenStudy (pphalke):

how do i do this? @benthicflare

OpenStudy (abdullahm):

All you need to do is the multiplication

OpenStudy (benthicflare):

@pphalke

OpenStudy (pphalke):

that was 3 days ago when i pulled the all nighter